OpenStudy (anonymous):

PLEASE HELP I WILL MEDAL AND FAN Solve for r: C = 2pir. r = C over 2 pi r = 2 c all over pi r = C over pi r = pi times C all over 2

OpenStudy (michele_laino):

Hint: if we divide both sides by 2*pi, we get: \[\Large \frac{C}{{2\pi }} = \frac{{2\pi r}}{{2\pi }}\] please simplify the right side

OpenStudy (anonymous):

2 pi r=c?

OpenStudy (usukidoll):

whenever there is a question like solve for r ... we need r to be by itself... so what do we need to do to make that happen?

OpenStudy (anonymous):

divide by 2

OpenStudy (usukidoll):

yes... we can do that... but there's more..

OpenStudy (anonymous):

2 pi

OpenStudy (usukidoll):

yes! so divide 2pi on both sides

OpenStudy (michele_laino):

hint: